# RG-Invariant Symmetry Ratio for QCD: A Study of $U(1)_{A}$ and Chiral Symmetry Restoration

### v1

@@ 6.5 RG interpretation and comparison with Pisarski–Wilczek analyses
The PW scenario applies strictly to the chiral limit $(m_{q}=0)$ at the exact infrared critical fixed point and therefore characterizes universality classes, rather than specific observables at finite quark mass. In physical QCD with $m_{q}\neq 0$, where the transition is a smooth crossover and no true critical fixed point exists, axial symmetry restoration must be defined operationally. In this work, we define “effective restoration” through the vanishing of the $\kappa_{AB}$ ratio within statistical uncertainties (Sec. 2.3). This condition is substantially weaker than the RG irrelevance criterion of PW: it does not imply enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or a change in universality class, but only that symmetry-breaking effects are not resolved within the present observables and statistical precision.
@@
Stage 2 (singlet sector): At a significantly higher temperature $T_{c}^{\mathrm{s}}\gg T_{c}^{\mathrm{ns}}$, the singlet ratio $\kappa_{TX}^{s}$ and the topological susceptibility $\chi_{t}$ also become consistent with zero, extending operational effective restoration to both singlet and nonsinglet sectors. This defines a more complete restoration scale than the PW critical temperature, which is determined solely by chiral symmetry in the massless limit and does not require observable axial degeneracy. Nevertheless, even at $T_{c}^{\mathrm{s}}$, this operational restoration does not by itself establish enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or determine the universality class, as explicit symmetry-breaking effects from finite quark masses and the anomaly remain present in the Lagrangian. Rather, $T_{c}^{\mathrm{s}}$ marks the temperature at which symmetry breaking becomes unresolvable in all measured channels, representing a stronger phenomenological restoration scale beyond the strict RG definition of criticality.
The renormalization-group analysis of Pelissetto-Vicari (PV) Pelissetto:2013hqa showed that for $N_{f}=2$ the $U(1)_{A}$-breaking operator is marginally irrelevant at the $O(4)$ fixed point, implying approximate $O(4)$ scaling in the chiral limit near the Pisarski–Wilczek critical temperature. Our results differ qualitatively from this scenario. At $T_{c}^{\mathrm{ns}}\sim 156$ MeV, we observe degeneracy in nonsinglet channels while singlet observables and the topological susceptibility remain nonvanishing, revealing a two-stage hierarchy of operational restoration. This structure does not follow from the fixed-point analysis, which applies strictly in the chiral limit and does not distinguish connected and disconnected contributions. Moreover, our temperatures $T_{c}^{\mathrm{ns}}$ and $T_{c}^{\mathrm{s}}$ are crossover scales in physical QCD rather than critical points, and the persistence of $\chi_{t}>0$ reflects the continued contribution of disconnected diagrams and the mass-dependent amplification of axial-anomaly effects. Thus, while marginal irrelevance provides an important conceptual backdrop, the hierarchical restoration observed here represents a distinct finite-mass phenomenon beyond the scope of the chiral-limit RG framework.
@@
Within this framework, our $\kappa_{AB}$ observables provide the first continuum-extrapolated, quantitative evidence for Stage 1, namely the operational restoration of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ in the nonsinglet (connected) sector near $T_{c}$. Establishing Stage 2, corresponding to full operational restoration including singlet channels, requires both $\kappa_{TX}^{s}\to 0$ and $\chi_{t}\to 0$ within uncertainties. Existing lattice studies show that topological fluctuations persist up to $T\gtrsim 2T_{c}$ (see e.g., Petreczky:2016vrs ; Borsanyi:2016ksw ; Chen:2022fid ; Athenodorou:2022aay ; Kotov:2025ilm ), indicating that anomaly effects remain active well above the crossover. However, a definitive continuum-limit determination combining chirally symmetric fermions and overlap-based topology measurements for both sea and valence sectors is still lacking. Such calculations are necessary to determine whether and at what temperature full effective restoration occurs.
@@ 6.6 Concluding remarks
Combined with future continuum-extrapolated determinations of tensor singlet channels and overlap fermions based measurements of $\chi_{t}$, this framework offers a comprehensive strategy for mapping the complete restoration pattern of chiral and axial symmetries in QCD.
@@ Appendix A Meson Operator Notation
$\rho^{a}$ (tensor) $\bar{q}\gamma_{4}\gamma_{i}t^{a}q$ Tensor ($T_{i}$) Nonsinglet ($ns$) $\chi_{\rho,T}=\chi_{T_{i}}^{ns}$
$\omega$ $\bar{q}\gamma_{4}\gamma_{i}q$ Tensor ($T_{i}$) Singlet ($s$) $\chi_{\omega}=\chi_{T_{i}}^{s}$
$h_{1}$ $\bar{q}\gamma_{5}\gamma_{4}\gamma_{i}q$ Axial-tensor ($X_{i}$) Singlet ($s$) $\chi_{h_{1}}=\chi_{X_{i}}^{s}$
@@
  • Spatial index $i=1,2,3$ (or $k=1,2,3$) is implicit for vector/tensor channels.
  • Singlet operators ($\sigma$, $\eta$, $\omega$, $h_{1}$) involve disconnected diagrams and require careful subtraction of power divergences.
@@
  • The tensor representation $\rho^{a}$ (via $T_{i}$) and $b_{1}^{a}$ (via $X_{i}$) probes $U(1)_{A}$ through a different Dirac structure than the scalar–pseudoscalar pair ($\delta$$\pi$).

### v2

@@ 6.5 RG interpretation and comparison with Pisarski–Wilczek analyses
The PW scenario applies strictly to the chiral limit $(m_{q}=0)$ at the exact infrared critical fixed point and therefore characterizes universality classes, rather than specific observables at finite quark mass. In physical QCD with $m_{q}\neq 0$, where the transition is a smooth crossover and no true critical fixed point exists, axial symmetry restoration must be defined operationally. In this work, we define “effective restoration” through the vanishing of the $\kappa_{AB}$ ratio within statistical uncertainties (section 2.3). This condition is substantially weaker than the RG irrelevance criterion of PW: it does not imply enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or a change in universality class, but only that symmetry-breaking effects are not resolved within the present observables and statistical precision.
@@
Stage 2 (singlet sector): At a significantly higher temperature $T_{c}^{\mathrm{s}}\gg T_{c}^{\mathrm{ns}}$, the singlet ratio $\kappa_{\omega_{T},h_{1T}}$ and the topological susceptibility $\chi_{t}$ also become consistent with zero, extending operational effective restoration to both singlet and nonsinglet sectors. This defines a more complete restoration scale than the PW critical temperature, which is determined solely by chiral symmetry in the massless limit and does not require observable axial degeneracy. Nevertheless, even at $T_{c}^{\mathrm{s}}$, this operational restoration does not by itself establish enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or determine the universality class, as explicit symmetry-breaking effects from finite quark masses and the anomaly remain present in the Lagrangian. Rather, $T_{c}^{\mathrm{s}}$ marks the temperature at which symmetry breaking becomes unresolvable in all measured channels, representing a stronger phenomenological restoration scale beyond the strict RG definition of criticality.
Pisarski and Wilczek (PW) Pisarski:1983ms identified two scenarios depending on whether the $U(1)_{A}$-breaking operator is relevant or irrelevant at the $O(4)$ fixed point, with the latter implying enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ and a fluctuation-induced first-order transition. The subsequent renormalization-group analysis of Pelissetto and Vicari (PV) Pelissetto:2013hqa established that for $N_{f}=2$ this operator is marginally irrelevant at the $O(4)$ fixed point, implying approximate $O(4)$ scaling in the chiral limit near the PW critical temperature. Our results differ qualitatively from this scenario. At $T_{c}^{\mathrm{ns}}\sim 156$ MeV, we observe degeneracy in nonsinglet channels while singlet observables and the topological susceptibility remain nonvanishing, revealing a two-stage hierarchy of operational restoration. This structure does not follow from the fixed-point analysis, which applies strictly in the chiral limit and does not distinguish connected and disconnected contributions. Moreover, our temperatures $T_{c}^{\mathrm{ns}}$ and $T_{c}^{\mathrm{s}}$ are crossover scales in physical QCD rather than critical points, and the persistence of $\chi_{t}>0$ reflects the continued contribution of disconnected diagrams and the mass-dependent amplification of axial-anomaly effects. Thus, while marginal irrelevance provides an important conceptual backdrop, the hierarchical restoration observed here represents a distinct finite-mass phenomenon beyond the scope of the chiral-limit RG framework.
@@
Within this framework, our $\kappa_{AB}$ observables provide the first continuum-extrapolated, quantitative evidence for Stage 1, namely the operational restoration of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ in the nonsinglet (connected) sector near $T_{c}$. It is worth noting that other ratios such as $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$, which involve singlet channels, are also formally RG-invariant, as the underlying renormalization constants satisfy $Z_{S}^{ns}=Z_{P}^{ns}=Z_{S}^{s}=Z_{P}^{s}$ (see section 2.4). However, in practice, even after contact term and vacuum subtractions, the singlet susceptibilities $\chi_{\sigma}$ and $\chi_{\eta}$ retain finite mixing with the topological susceptibility $\chi_{t}$, which must be disentangled to obtain fully renormalized quantities. This additional complication, together with the noise from disconnected diagrams, makes these ratios less accessible in the present study.
@@
Establishing Stage 2, corresponding to full operational restoration including singlet channels, requires both $\kappa_{\omega_{T},h_{1T}}\to 0$ and $\chi_{t}\to 0$ within uncertainties. Existing lattice studies show that topological fluctuations persist up to $T\gtrsim 2T_{c}$ (see e.g., Petreczky:2016vrs ; Borsanyi:2016ksw ; Chen:2022fid ; Athenodorou:2022aay ; Kotov:2025ilm ), indicating that anomaly effects remain active well above the crossover. However, a definitive continuum-limit determination combining chirally symmetric fermions and overlap-based topology measurements for both sea and valence sectors is still lacking. Such calculations are necessary to determine whether and at what temperature full effective restoration occurs.
@@ 6.6 Concluding remarks
Combined with future continuum-extrapolated determinations of tensor singlet channels and overlap-fermion-based measurements of $\chi_{t}$, this framework offers a comprehensive strategy for mapping the complete restoration pattern of chiral and axial symmetries in QCD.
@@ Appendix A Meson Operator Notation
$\omega$ $\bar{q}\gamma_{i}q$ Vector ($V_{i}$) Singlet ($s$) $\chi_{\omega}=\chi_{V_{i}}^{s}$
$h_{1}$ $\bar{q}\gamma_{5}\gamma_{i}q$ Axial-vector ($A_{i}$) Singlet ($s$) $\chi_{h_{1}}=\chi_{A_{i}}^{s}$
$\rho_{T}^{a}$ $\bar{q}\gamma_{4}\gamma_{i}t^{a}q$ Tensor ($T_{i}$) Nonsinglet ($ns$) $\chi_{\rho_{T}}=\chi_{T_{i}}^{ns}$
@@
$\omega_{T}$ $\bar{q}\gamma_{4}\gamma_{i}q$ Tensor ($T_{i}$) Singlet ($s$) $\chi_{\omega_{T}}=\chi_{T_{i}}^{s}$
$h_{1T}$ $\bar{q}\gamma_{5}\gamma_{4}\gamma_{i}q$ Axial-tensor ($X_{i}$) Singlet ($s$) $\chi_{h_{1T}}=\chi_{X_{i}}^{s}$
@@
  • Spatial index $i=1,2,3$ (or $k=1,2,3$) is implicit for vector and tensor channels.
  • Singlet operators ($\sigma$, $\eta$, $\omega$, $h_{1}$, $\omega_{T}$, $h_{1T}$) involve disconnected diagrams and require careful subtraction of power divergences.
@@
  • The tensor representations $\rho_{T}^{a}$ (via $T_{i}$) and $b_{1}^{a}$ (via $X_{i}$) probe $U(1)_{A}$ through a different Dirac structure than the scalar–pseudoscalar pair ($\delta$$\pi$).
  • The vector singlet $\omega$ and axial-vector singlet $h_{1}$ are the $I=0$ partners under $U(1)_{A}$, while the tensor singlet $\omega_{T}$ and axial-tensor singlet $h_{1T}$ play analogous roles in the tensor channel.